Twice a first number plus a second number is 18. The first number is five less than three times the second number. Find the numbers. Show equations.
Answer by nyc_function(2645)

2x + y = 18
x = 3y - 5
Those are your equations.
Can you finish now?
=====================================================
I got your e-mail. Here is the entire problem solved.

Good morning. I got your reply lasy night. Here is how we do your math problem.
Twice a first number plus a second number is 18. The first number is five less than three times the second number. Find the numbers. Show equations.


Solution:


Let x = one number
let y = the other number


2x + y = 18
x = 3y - 5


I will use the substitution method, which is the easiest to follow.


Plug 3y - 5 into the other equation and solve for y.


2(3y - 5) + y = 18


6y - 10 + y = 18


7y = 18 + 10


7y = 28


y = 28/7


y = 4


Since we now know y = 4, we plug 4 for y into EITHER of the original equations created from the word problem.


I will use x = 3y - 5.


Let y = 4


x = 3(4) - 5


x = 12 - 5


x = 7


The numbers are 4 and 7.


Did you follow?


Guido